OpenStudy (anonymous):

A circle is tangent to the x axis at x=5 and has one y intercept at y=3, determine the other y intercept and the equation of the circle

OpenStudy (dumbcow):

 from the picture, you see the center is up "k" units from x-axis but that is also the radius of the circle --> k = r general equation for circle \[(x-h)^{2} + (y-k)^{2} = r^{2}\] also note that h=5 since the center is directly above point (5,0) Now to solve for r , plug in other given point (0,3) for x,y \[(0-5)^{2}+(3-r)^{2} = r^{2}\] \[25 + (9-6r+r^{2}) = r^{2}\] \[r = \frac{34}{6} = \frac{17}{3}\] circle equation: \[(x-5)^{2} + (y-17/3)^{2} = (17/3)^{2}\]

OpenStudy (dumbcow):

to find other y_intercept, plug in x=0 and solve for y